# Cosmic-Ray Feedback from Supernovae in a Stratified Interstellar Medium

### v1

@@ Abstract
Each supernova’s energy drives interstellar medium (ISM) turbulence and can help launch galactic winds. What difference does it make if $10\%$ of the energy is initially deposited into cosmic rays? To answer this question and study cosmic-ray feedback, we perform galactic patch simulations of a stratified ISM. We compare two magnetohydrodynamic and cosmic ray (MHD+CR) simulations, which are identical except for how each supernova’s energy is injected. In one, $10\%$ of the energy is injected as cosmic-ray energy and the rest is thermal. In the other case, energy injection is strictly thermal. We find that cosmic-ray injections (1) drive a faster vertical motion with more mass, (2) produce a more vertically oriented magnetic field, and (3) increase the scale height of warm gas outside the midplane $(z\gtrsim 0.5\,\mathrm{kpc})$. Both simulations show the formation of cold clouds (with a total mass fraction $>50\%$) through the Parker instability and thermal instability. We also show that the Parker instability leads to a decorrelation of cosmic-ray pressure and gas density. Finally, our simulations show that a vertical magnetic field can lead to a significant decrease in the calorimetric fraction for injected cosmic rays.
@@ 1 Introduction
A tenth, $10\%$, or $0.1$ of anything is a small fraction. Neglecting such a small component is a reasonable assumption in many situations. An example of this small component is cosmic rays in the interstellar medium (ISM): $10\%$ of each supernova’s kinetic energy is all that is needed to explain the amount of these high-energy particles in our galaxy’s ISM (Baade & Zwicky, 1934; Blandford & Ostriker, 1978). More recent work has illustrated that the fraction produced in a shock could be as high as $20\%$ (Caprioli & Spitkovsky, 2014). That fraction could still be considered negligible, especially on the long time scales $(t\gtrsim 100\,\mathrm{Myr})$ and long length scales $(L\gtrsim 1\,\mathrm{kpc})$ of galaxy evolution, since supernovae drive turbulence and heat gas at small scales $(L\sim 1\mathrm{pc})$.
@@
However, other work has shown that cosmic rays are not negligible on large scales (see recent reviews Zweibel (2017); Ruszkowski & Pfrommer (2023); Owen et al. (2023)). They are in approximate pressure/energy equipartition with the thermal gas and magnetic field (Ferrière, 2001). Gradients in cosmic-ray pressure can drive outflows or fountain flows from a galactic disk, removing gas that could otherwise form stars (Chan et al., 2022; Tsung et al., 2023; Armillotta et al., 2024; Thomas et al., 2024). Bottlenecks created by cosmic-ray streaming can destroy and disrupt cold gas clouds (Wiener et al., 2019; Bustard & Zweibel, 2021; Tsung et al., 2022). Cosmic-ray transport is faster parallel to the magnetic field, and this anisotropic transport can impact the turbulent energy cascade of the thermal gas (Habegger et al., 2024).
@@
These feedback mechanisms are stifled if cosmic rays are transported out of the galaxy too quickly. In order to keep global galaxy simulations (cosmological zoom-in simulations) in agreement with observed $\gamma$-ray luminosities, a cosmic-ray diffusion coefficient of about one order of magnitude higher than the standard value of $3\times 10^{28}\,\mathrm{cm}^{2}\,\mathrm{s}^{-1}$ derived from the B/C ratio (Jones et al., 2001; Evoli et al., 2020) is often invoked, following the simulations of Chan et al. (2019). This diffusion coefficient amplification means that the cosmic rays escape the galaxy before they ever affect the dynamics of the thermal gas. This result brings us back to the idea that cosmic rays produce only a marginal effect; including them just makes stellar feedback slightly more efficient.
@@
In this work, we examine how individual cosmic-ray sources eventually adjust the dynamics and vertical structure of the multiphase ISM, becoming a significant component of stellar feedback. We previously examined how a single localized cosmic-ray energy injection on a magnetic flux tube could disrupt a stratified ISM (Habegger et al., 2023). This work builds on that study to consider many injections throughout the galactic midplane. Our simulations evolve a slab of stratified ISM, with midplane parameters that match the solar neighborhood. After a significant fountain flow develops in the first $200\,\mathrm{Myr}$, with magnetic field geometry similar to the Parker instability (Parker, 1966), the slab reaches a steady state (lasting $\gtrsim 100\,\mathrm{Myr}$). During the transition to the steady state, the cosmic rays are able to escape quickly without interacting with the thermal gas. In the steady state, we show how individual cosmic-ray injections allow cosmic-ray pressure to be the primary force outside the galactic midplane.
@@
Figure 1 is a snapshot of one of our simulations, which illustrates the complex temperature and magnetic field structures that form over time in the stratified, multiphase, and magnetized ISM. The temperature varies over several orders of magnitude. Individual magnetic field lines stretch over $1\,\mathrm{kpc}$ in length. Considering the full scale of a galaxy, those magnetic field lines would connect the galactic plane and the galactic halo. We find that these field lines act as escape highways for injected cosmic rays, keeping the simulation below the calorimetric limit. These extended field lines result from expanding plumes that push the horizontal magnetic field out of the galactic plane. Those plumes drive the creation of cold gas in valleys, the same structure predicted for the Parker instability (Parker, 1966). These valleys result in a decorrelation of cosmic-ray pressure and gas density. The decorrelation is visible in Figure 1, where the cold gas (blue contours) and high cosmic-ray pressure regions (brown contours) alternate in the $y$ direction, which is perpendicular to the mean magnetic field.
Figure 1: 3D snapshot at $t=180\,\mathrm{Myr}$ from our CRInj simulation (see Table 1). Small blue structures near $z=0$ are contours around cold gas $(T\leq 500\,\mathrm{K})$, and the large red sheets at $|z|>1000\,\mathrm{pc}$ are contours around hot gas $(T\geq 2\times 10^{4}\,\mathrm{K})$. Green lines trace the magnetic field, and darker green corresponds to stronger magnetic field. The brown contours in the $z=0$ plane surround regions of high cosmic-ray pressure.
@@ 2 Methods
These equations describe the evolution of a magnetized two-fluid system. The first fluid, thermal gas, has a mass density $\rho$, velocity $\mathbf{u}$, pressure $P_{g}$. The second fluid, made up of cosmic rays (i.e. a relativistic, non-thermal fluid), is described by a total energy density $E_{c}$ and flux $\mathbf{F}_{c}$. The cosmic-ray transport coefficient $\sigma_{c}$ is a diagonal matrix which governs how cosmic-ray energy moves in the simulation, $\mathbf{v}_{s}$ is the streaming velocity, and the modified speed of light $V_{m}$ sets the maximum transport speed. We evolve the magnetic field $\mathbf{B}$ from Faraday’s Law in the ideal MHD approximation (Eqn. 2). The remaining variables are the total magnetohydrodynamic energy density of the first fluid $E=P_{g}/(\gamma-1)+\rho u^{2}/2+B^{2}/2$, the gravitational field $\mathbf{g}$, and the radiative heating and cooling function $\mathcal{L}(n,T)$, which depends on the number density ($n=\rho/m$, where $m$ is particle mass) and temperature $(T=P_{g}m/(\rho k_{B}))$ of the thermal gas.
For the cosmic-ray transport, we allow the cosmic rays to diffuse along the local magnetic field at the canonical Milky Way value of $\kappa_{\parallel}=3\times 10^{28}\,\mathrm{cm}^{2}\mathrm{s}^{-1}$ (Jones et al. (2001), for more recent observations see Evoli et al. (2019, 2020)). This anisotropic transport is created by splitting the transport coefficient into components perpendicular and parallel to the local magnetic field direction $\hat{b}$ in each cell (see Jiang & Oh (2018)):
@@
We also include the effects of cosmic-ray streaming transport. The additional heating from dissipation of Alfven waves is included in the energy equations (MHD energy in Equation 4, cosmic-ray energy in Equation 5). The additional transport appears in the transport matrix by including the following term:
@@
Additionally, it is important to note that the streaming velocity should be
$\mathbf{v}_{s}=-\frac{\mathbf{B}}{\sqrt{4\pi\rho}}\frac{\mathbf{B}\cdot\gradient{P_{c}}}{\left|\mathbf{B}\cdot\gradient{P_{c}}\right|}$ (9)
for the equations in Jiang & Oh (2018) to match with other formulations of the cosmic-ray energy equation (e.g. McKenzie & Voelk (1982); Breitschwerdt et al. (1991); Zweibel (2017)).
@@
In the simulations used for this work, we used $V_{m}=0.1c$. See Appendix C of Habegger et al. (2023) for an extended $V_{m}$ convergence study. Our largest flow velocities are on the order of $100\,\mathrm{km}\,\mathrm{s}^{-1}$, much lower than our chosen $V_{m}$.
Parameter CRInj NoCRInj
@@
$E_{\mathrm{inj,cr}}/\mathrm{erg}$ $0.1\cdot 10^{51}$ 00
@@ 2.1 Setup & Initial Conditions
We present two simulations, with identical initial conditions, physics, and random seed. The only difference is how supernova energy is injected. In one simulation, named CRInj, each supernova has $10\%$ of its energy injected as cosmic-ray energy instead of thermal energy. In the second simulation, named NoCRInj, each supernova has all of its energy injected as thermal energy, with no cosmic-ray component. The important parameters for each simulation are show in Table 1.
The simulations are on a Cartesian grid, with the $\hat{x}$ and $\hat{y}$ directions spanning the midplane of the slab and the $\hat{z}$ direction extending perpendicular to the midplane. The midplane has an area $1\,\mathrm{kpc}\times 1\,\mathrm{kpc}$, but the simulation extends ±$\pm 2.4\,\mathrm{kpc}$, above and below the midplane, for a total simulation volume of $1\,\mathrm{kpc}^{2}\times 4.8\,\mathrm{kpc}$. Each direction has a resolution of $\Delta x=\Delta y=\Delta z=10\,\mathrm{pc}$, resulting in $100$ resolution elements in the $\hat{x}$ and $\hat{y}$ directions, but $480$ in the $\hat{z}$ direction.
@@
$\mathbf{g}(z)=-2\pi G\Sigma_{*}\hat{z}\tanh\left(\frac{z}{H_{*}}\right).$ (12)
@@
All velocities are initially set to zero, and the $\hat{z}$ and $\hat{y}$ components of the magnetic field are set to 00 everywhere.
[…]
The finite size of the galaxy results in less acceleration because there are less stars than assumed in the infinite plane approximation, producing less acceleration. The dark matter’s effect is less obvious: the dark matter halo produces a radial acceleration, towards the galactic center. This additional matter does mean a larger acceleration, but it will be in the radial direction, not only in the vertical direction. This acceleration is still smaller than what we get assuming an infinite plane.
So, we are over-estimating the vertical gravitational acceleration on gas at large $z$. Any outflow velocities we find in Section 3 are therefore weaker than in a more realistic gravitational acceleration profile. Another consideration would be the Coriolis force, but that would again only serve to increase the velocities beyond what we find in our simulation.
@@ 2.2 Heating and Radiative Cooling
We use the heating rate from Inoue et al. (2006), a constant $\Gamma=2\cdot 10^{-26}\,\mathrm{erg}\,\mathrm{s}^{-1}$ from photoelectric heating by dust grains (Inoue et al., 2006; Draine, 2011). With all the parameters of Equation 17 set, we have an equilibrium curve defined by $\Gamma=n\Lambda(T)$. We show this curve in pressure-density phase space as a solid black line in the top plot of Figure 2. As mentioned in Section 2.1, we start with an isothermal atmosphere. The number density ranges from $n=1$ to approximately $n\approx 10^{-4}$. This isothermal atmosphere quickly $(\lesssim 10\,\mathrm{Myr})$ relaxes to the equilibrium curve in that density range before the energy injections significantly adjust the structure.
@@
In Figure 2, the red areas are where $\tau>0$ in Equation 20, giving net heating. The blue areas are where $\tau<0$, resulting in net cooling. We plot the isochoric time because we use an operator split approach, including the heating and radiative cooling as an isochoric process. Adiabatic compression and expansion are already included in the energy evolution (see Equation 4). This treatment is common, e.g. see Townsend (2009). Unlike the exact method in Townsend (2009), we use an explicit integration for the heating and radiative cooling, with order set by the MHD time integrator. Explicit integration is reasonable for our simulations because our cooling time is longer than the simulation’s time step $(\Delta t=\mathrm{CFL}\times\Delta x/V_{m}\sim 65.2\,\mathrm{yr})$, which is set by the modified speed of light $V_{m}$.
Examining Figure 2, we see at higher densities $n\gtrsim 10^{-2}\,\mathrm{cm}^{-3}$ the gas cools or heats on short timescales, until it again reaches the equilibrium curve where heating and cooling balance. We also plot our floor temperature of $T=30\,\mathrm{K}$ as a dashed line in the top plot of Figure 2. Whenever the gas reaches the temperature floor, we turn off the radiative heating and cooling terms. If the gas is compressed or expanded by gas motions and brought above the temperature floor, then the terms are again included.
An important caveat is that our resolution of $10\,\mathrm{pc}$ is rather large when attempting to study the thermal instability created by the cooling curve from Inoue et al. (2006). Therefore, the shapes of cold dense structures formed in our simulation are not reliable. There should be fragmentation down to smaller scales than we resolve. The resolution also limits the peak gas density in the simulations. However, we will see in Section 3 that the development of the thermal instability is resolved by multiple gas cells. Without a curve like the one from Inoue et al. (2006), we would not form cold gas through Field’s thermal instability (Field, 1965) because there would not be a strong turnover in the pressure-density phase space. By using the Inoue et al. (2006) curve, the initial production of unstable gas is closer to what is expected in the Milky Way’s ISM.
Another aspect of our implementation to note is that our radiative heating mechanism of a constant $\Gamma$ for all gas densities and vertical positions is unrealistic. The heating rate should fall off with vertical position $z$, allowing more radiative cooling to occur outside the midplane. This choice makes it harder for gas to cool at large $z$ where gas densities are also lower (lower density reduces the cooling term more than it reduces the heating term in Equation 17). However, choosing a constant $\Gamma$ makes the thermodynamic equilibrium curve (top panel of Figure 2) the same for the entire simulation volume, instead of the curve being dependent on position. This simplification makes it easier isolate the effect of the energy injections on the state of the gas.

### v2

# Cosmic-Ray Feedback from Supernovae in a Parker-Unstable Medium

@@ Abstract
Supernova energy drives interstellar medium (ISM) turbulence and can help launch galactic winds. What difference does it make if $10\%$ of the energy is initially deposited into cosmic rays? To answer this question and study cosmic-ray feedback, we perform galactic patch simulations of a stratified ISM. We compare two magnetohydrodynamic and cosmic ray (MHD+CR) simulations, which are identical except for how each supernova’s energy is injected. In one, $10\%$ of the energy is injected as cosmic-ray energy. In the other case, energy injection is strictly thermal and kinetic. We find that cosmic-ray injections drive a faster, hotter, and more massive outflow long after the injections occur. Both simulations show the formation of cold clouds (with a total mass fraction $>50\%$) through the Parker instability and thermal instability. The Parker instability simultaneously produces high mass loading factors $\eta>10^{3}$ as it requires few supernovae. We also show how the Parker instability naturally leads to a decorrelation of cosmic-ray pressure and gas density. This decorrelation leads to a significant decrease in the calorimetric fraction for injected cosmic rays, but it depends on having a highly resolved magnetic field.
@@ 1 Introduction
Figure 1: 3D snapshots of each simulation at $t=200\,\mathrm{Myr}$ (see Table 1). The top row shows volume renders from our CRInj simulation, and the bottom row shows renders from our CRBkg simulation. The left side figures show a larger scale picture illustrating the hot wind and exgtended magnetic field, whereas the right side figures show a zoom-in of the midplane region. The blue structures are contours around cold gas $(T\leq 500\,\mathrm{K})$, and the red contours are around hot gas $(T\geq 5\times 10^{4}\,\mathrm{K})$. Green lines trace the magnetic field, and darker green corresponds to stronger magnetic field (see colorbar in left column figures). The orange contours in the $z=0$ surround regions of high cosmic-ray pressure. The red blobs in the midplane show expanding supernova. In the CRInj simulation, these supernovae are encased by orange contours showing where they dumped cosmic ray energy along the local magnetic field.
@@
A tenth, $10\%$, or $0.1$ of anything is a small fraction. Neglecting such a small component is a reasonable assumption in many situations. An example of this small component is cosmic rays in the interstellar medium (ISM): $10\%$ of each supernova’s energy is all that is needed to explain the amount of these high-energy particles in our galaxy’s ISM (Baade & Zwicky, 1934; Blandford & Ostriker, 1978). More recent work has illustrated that the fraction produced in a shock could be as high as $20\%$ (Caprioli & Spitkovsky, 2014). That fraction could still be considered negligible, especially on the long time scales $(t\gtrsim 100\,\mathrm{Myr})$ and long length scales $(L\gtrsim 1\,\mathrm{kpc})$ of galaxy evolution, since supernovae drive turbulence and heat gas at small scales $(L\sim 10\mathrm{pc})$.
@@
However, other work has shown that cosmic rays are not negligible on large scales (see recent reviews Zweibel (2017); Ruszkowski & Pfrommer (2023); Owen et al. (2023)). They are in approximate pressure/energy equipartition with the thermal gas and magnetic field (Ferrière, 2001). Gradients in cosmic-ray pressure can drive outflows or fountain flows from a galactic disk, removing gas that could otherwise form stars (Chan et al., 2022; Tsung et al., 2023; Armillotta et al., 2024; Thomas et al., 2024). Bottlenecks created by cosmic-ray streaming can destroy and disrupt cold gas clouds (Wiener et al., 2017, 2019; Bustard & Zweibel, 2021; Tsung et al., 2022). Cosmic-ray transport is faster parallel to the magnetic field, and this anisotropic transport can impact the turbulent energy cascade of the thermal gas (Habegger et al., 2024).
@@
These feedback mechanisms are stifled if cosmic rays are transported out of the galaxy too quickly. In order to keep global galaxy simulations (and cosmological zoom-in simulations) in agreement with observed $\gamma$-ray luminosities, a cosmic-ray diffusion coefficient of about one order of magnitude higher than the standard value of $3\times 10^{28}\,\mathrm{cm}^{2}\,\mathrm{s}^{-1}$ derived from the B/C ratio (Jones et al., 2001; Evoli et al., 2020) is often invoked, following the simulations of Chan et al. (2019). This diffusion coefficient amplification means that the cosmic rays escape the galaxy before they ever affect the dynamics of the thermal gas. This result brings us back to the idea that cosmic rays produce only a marginal effect; including them just makes stellar feedback slightly more efficient.
@@
In this work, we examine how individual cosmic-ray sources eventually adjust the dynamics and vertical structure of the multiphase ISM, becoming a significant component of stellar feedback. We previously examined how a single localized cosmic-ray energy injection on a magnetic flux tube could disrupt a stratified ISM (Habegger et al., 2023). This work builds on that study to consider multiple injections throughout the galactic midplane. Our simulations evolve a slab of stratified ISM, with midplane parameters that match the solar neighborhood. After a significant fountain flow develops in the first $200\,\mathrm{Myr}$, with magnetic field geometry similar to the Parker instability (Parker, 1966), the slab reaches a steady state (lasting $\gtrsim 100\,\mathrm{Myr}$). During the transition to the steady state, the cosmic rays are able to escape quickly without interacting with the thermal gas. This escape significantly decreases the $\gamma$-ray luminosity.
@@
Figure 1 is a set of volume renders from our simulations at a time of $t=200\,\mathrm{Myr}$, which illustrate the complex temperature and magnetic field structures that form over time in the stratified, multiphase, and magnetized ISM. The temperature varies over several orders of magnitude (blue contours surround gas with $T<500\,\mathrm{K}$ and red contours surround gas with $T>5\cdot 10^{4}\,\mathrm{K}$). Regions of high cosmic-ray pressure (orange contours) follow the magnetic field (in green), inflating magnetic flux tubes. Some magnetic field lines stretch perpendicular to the disk, over $1\,\mathrm{kpc}$ in length, despite initially only being directed horizontally. Considering the full scale of a galaxy, those magnetic field lines would connect the galactic plane and the galactic halo. We find that these field lines act as escape highways for injected cosmic rays, taking the simulation below the calorimetric limit. These extended field lines result from expanding plumes that push the horizontal magnetic field out of the galactic plane. Those plumes drive the creation of cold gas in valleys, the same structure predicted for the Parker instability (Parker, 1966). These valleys result in a decorrelation of cosmic-ray pressure and gas density. The decorrelation is visible in Figure 1, where the cold gas (blue contours) and high cosmic-ray pressure regions (orange contours) alternate in the direction perpendicular to the mean magnetic field. While the Parker instability drives this decorrelation in our simulations, the general process of accreting cold gas from the galactic halo should have the same effect as long as the magnetic field is well-resolved.
@@ 2 Methods
These equations describe the evolution of a magnetized two-fluid system. The first fluid, thermal gas, has a mass density $\rho$, velocity $\mathbf{u}$, pressure $P_{g}$. The second fluid, made up of cosmic rays (i.e. a relativistic, non-thermal fluid), is described by a total energy density $E_{c}$ and flux $\mathbf{F}_{c}$. The cosmic-ray transport coefficient $\sigma_{c}$ is a diagonal matrix which governs how cosmic-ray energy moves in the simulation, $\mathbf{v}_{s}$ is the streaming velocity, and the modified speed of light $V_{m}$ sets the maximum transport speed. We evolve the magnetic field $\mathbf{B}$ from Faraday’s Law in the ideal MHD approximation (Eqn. 2). The remaining variables are the total magnetohydrodynamic energy density of the first fluid $E=P_{g}/(\gamma-1)+\rho u^{2}/2+B^{2}/2$, the gravitational field $\mathbf{g}$, and the radiative heating and cooling function $\mathcal{L}(n,T)$, which depends on the number density ($n=\rho/m$, where $m$ is particle mass) and temperature $(T=P_{g}m/(\rho k_{B}))$ of the thermal gas. We assume hydrogen gas with a constant molecular weight of $\mu=1$, so the particle mass is $m=\mu m_{H}=m_{H}$.
For the cosmic-ray transport, we allow the cosmic rays to diffuse along the local magnetic field at the canonical Milky Way value of $\kappa_{\parallel}=3\times 10^{28}\,\mathrm{cm}^{2}\mathrm{s}^{-1}$ (Jones et al. (2001), for more recent determinations see Evoli et al. (2019, 2020)). This anisotropic transport is created by splitting the transport coefficient into components perpendicular and parallel to the local magnetic field direction $\hat{b}$ in each cell (see Jiang & Oh (2018)):
@@
We also include the effects of cosmic-ray streaming transport. Streaming at the local Alfven speed is dominant when cosmic rays generate the same Alfven waves that they gyro-resonantly scatter off (Zweibel, 2017). This process leads to additional heating because the Alfven waves dissipate after scattering the cosmic rays, taking energy from the cosmic rays into the thermal gas. The heating of gas (and decrease in cosmic ray energy) are proportional to the streaming speed $\mathbf{v}_{s}$ and appear in the energy equations (MHD energy in Equation 4, cosmic-ray energy in Equation 5). The additional transport appears in the transport matrix by including the following term:
@@
Additionally, it is important to note that the streaming velocity is the Alfven velocity directed down the cosmic ray pressure gradient (McKenzie & Voelk, 1982; Breitschwerdt et al., 1991; Zweibel, 2017; Jiang & Oh, 2018):
$\mathbf{v}_{s}=-\frac{\mathbf{B}}{\sqrt{4\pi\rho}}\frac{\mathbf{B}\cdot\gradient{P_{c}}}{\left|\mathbf{B}\cdot\gradient{P_{c}}\right|}.$ (9)
Note that we use the full gas density $\rho$ in Equation 9, not the ion density. In reality, the ion density should appear in the definition of the Alfven speed for plasma with an ionization fraction less than one. This simplification only impacts the cold gas which forms in our simulations.
In cold gas, the ionization fraction decreases so the gas density and ion density disagree. Additionally, Alfven waves are damped making all gyroresonant scattering of cosmic rays weaker. Other works have successfully implemented and studied the effect of temperature dependent cosmic ray transport (Farber et al., 2018; Bustard & Zweibel, 2021; Armillotta et al., 2024; Habegger et al., 2024).
However, we neglect this effect in these simulations because our chosen diffusion rate is high enough for the cold gas. Diffusion across an $100\mathrm{pc}$ cloud occurs at a speed $\kappa_{\parallel}/L=10^{3}\mathrm{km}\mathrm{s}^{-1}$. For smaller clouds, the rate of diffusion is even faster. Increasing the rate further by including more complete physics for cosmic ray transport would have a minimal impact on our results because diffusion already rapidly smooths cosmic ray pressure gradients in cold gas.
@@
In the simulations used for this work, we set $V_{m}=0.1c$. See Appendix C of Habegger et al. (2023) for an extended $V_{m}$ convergence study. Our largest flow velocities are on the order of $100\,\mathrm{km}\,\mathrm{s}^{-1}$, much lower than our chosen $V_{m}$.
Parameter CRInj CRBkg
@@
$E_{\mathrm{inj,cr}}/\mathrm{erg}$ $1.0\cdot 10^{51}$ 00
@@ 2.1 Setup & Initial Conditions
We present two simulations, with identical initial conditions. The only difference is how supernova energy is injected. In one simulation, named CRInj, each supernova has $10\%$ of its energy injected as cosmic-ray energy instead of thermal energy. In the second simulation, named CRBkg, no cosmic-ray energy is injected with each supernova. Instead, the only cosmic rays are from the initial background we put in to match the observed estimates that cosmic ray pressure is similar in value to the thermal gas pressure (i.e. $P_{c}\sim P_{g}$.) The important parameters for each simulation are shown in Table 1.
The simulations are on a Cartesian grid, with the $\hat{x}$ and $\hat{y}$ directions spanning the midplane of the slab and the $\hat{z}$ direction extending perpendicular to the midplane. The midplane has an area $1\,\mathrm{kpc}\times 1\,\mathrm{kpc}$, but the simulation extends ±$\pm 2.4\,\mathrm{kpc}$, above and below the midplane, for a total simulation volume of $1\,\mathrm{kpc}^{2}\times 4.8\,\mathrm{kpc}$. In the $x$ and $y$ directions, the boundary conditions are periodic. In the $z$ direction, we use a vacuum or ‘diode’ boundary condition which only allows outflow and no inflow.
Each direction has a resolution of $\Delta x=\Delta y=\Delta z=5\,\mathrm{pc}$, resulting in $200$ resolution elements in the $\hat{x}$ and $\hat{y}$ directions, but $960$ in the $\hat{z}$ direction. This resolution, combined with momentum injection (see section 2.3), is necessary for an accurate model of supernova feedback in a multiphase ISM (Kim & Ostriker, 2015). In particular, both are necessary for resolving the Sedov-Taylor phase and the transition to the snowplow phase (Draine, 2011).
The simulations are initially in hydrostatic equilibrium, and the setup is identical to the one in Habegger et al. (2023). We use a gravitational acceleration profile
$\mathbf{g}(z)=-2\pi G\Sigma_{*}\hat{z}\tanh\left(\frac{z}{H_{*}}\right),$ (12)
@@
All velocities are initially set to zero, and the $\hat{z}$ and $\hat{y}$ components of the magnetic field are set to zero everywhere.
@@ 2.2 Heating and Radiative Cooling
We use the heating rate from Inoue et al. (2006), a constant $\Gamma=2\cdot 10^{-26}\,\mathrm{erg}\,\mathrm{s}^{-1}$ from photoelectric heating by dust grains (Inoue et al., 2006; Draine, 2011). With all the parameters of Equation 17 set, we have an equilibrium curve defined by $\Gamma=n\Lambda(T)$. We show this curve in pressure-density phase space as a solid black line in the top plot of Figure 2. As mentioned in Section 2.1, we start with an isothermal atmosphere. The number density ranges from $n=1$ to approximately $n\approx 10^{-4}$. This isothermal atmosphere quickly $(\lesssim 1\,\mathrm{Myr})$ relaxes to the equilibrium curve in that density range before the energy injections significantly adjust the structure.
@@
In Figure 2, the red areas are where $\tau>0$ in Equation 20, giving net heating. The blue areas are where $\tau<0$, resulting in net cooling. We plot the isochoric time because we use an operator split approach, including the heating and radiative cooling as an isochoric process. Adiabatic compression and expansion are already included in the energy evolution (see Equation 4). This treatment is common, e.g. see Townsend (2009). Unlike the exact method in Townsend (2009), we use an explicit integration for the heating and radiative cooling, with order set by the MHD time integrator. Explicit integration is reasonable for our simulations because our cooling time is longer than the simulation’s time step $(\Delta t=\mathrm{CFL}\times\Delta x/V_{m}\sim 32.6\,\mathrm{yr})$, which is set by the modified speed of light $V_{m}$.
Examining Figure 2, we see at higher densities $n\gtrsim 10^{-2}\,\mathrm{cm}^{-3}$ the gas cools or heats on short timescales, until it again reaches the equilibrium curve where heating and cooling balance. We also plot our floor temperature of $T=30\,\mathrm{K}$ as a dashed line in the top plot of Figure 2. This floor is necessary to avoid numerical errors, as well as to stop the formation of molecular gas which we do not accurately model. Whenever the gas reaches the temperature floor, we turn off the radiative heating and cooling terms. If the gas is compressed or expanded by gas motions and brought above the temperature floor, then the terms are again included.